the measure of the largest angle of a triangle is 45 degrees moe than the measure of the smallest angle. The middle angle measures 65 degrees. What are the measures of the largest and smallest angles?
Answer by checkley77(12569)

x+45+65=180
x=180-45-65
x=70 degrees for the missing angle.
Proof:
70+45+65=180
180=180